OpenStudy (anonymous):

find the difference quotient of f(x) = (5) / (x + 9)

OpenStudy (anonymous):

\[\frac{\Delta F(p)}{\Delta p}=\frac{f(p+\Delta p)-f(p)}{\Delta p}\]

OpenStudy (anonymous):

hold on i thought tyler was going to do it sorry

OpenStudy (anonymous):

\[\frac{-5}{(x+9)^2+h(x+9)}\]

OpenStudy (anonymous):

i believe that is correct above cause if i wanted the derivative of your function as h -> 0 it would cancelt the h(x+9) which would get me the derivative

OpenStudy (anonymous):

do you have the answer and sorry about the wait

OpenStudy (anonymous):

yeah, but when I entered it, I was told the true answer is \[-5/(x+9)(x+h+9)\]

OpenStudy (anonymous):

I appreciate your help, I just needed to figure out how to break it down.